OpenStudy (anonymous):

Algebra II, please help! Triple equation system--v 2x + y - z = 7 -x - 2y + 4z = -15 x + 2y - 5z = 17 Which ordered triple is solution of this system? a. (2,6,5) b. (1,3,2) c. (2,5,2) d. (3,1,0)

OpenStudy (anonymous):

I know how to do this because I am actually in Algebra II as well. So do you know what the first step is to solving it?

OpenStudy (anonymous):

not really :/ can u walk me thru it ?

OpenStudy (anonymous):

Yes, I will be happy too. Okay so first thin first. You want to plug in the #'s. For example: 2*2 + 6 - 5 = 7. -2 - 2*6 + 4*5 = -15 2 + 2*6 - 5*5 = 17 So you will see if they actually = 7, -15, and 17. And if that doesn't work then plug in the next set and see if that matches.

OpenStudy (inkyvoyd):

that works, but if they don't give you the answers you are screwed.

OpenStudy (inkyvoyd):

oftentimes to prevent this the question asks you for x+y+z...

OpenStudy (inkyvoyd):

you can eliminate 2 variables with a simple look at the last two equations: -x - 2y + 4z = -15 (2) x + 2y - 5z = 17 (3) simply add the two 0+0-z=2

OpenStudy (inkyvoyd):

simply solve for z, and you have eliminated a variable.

OpenStudy (anonymous):

you can always add together two equations. if one equation has the same constant positive in the one and negative in the other, it is "eliminated": example: ´´´´´´´´ a + 2 b = 9 -a + b = 2 ---------- 3 b = 11 variable "a" was "eliminated" by adding the example equations. the same is possible in your problem, see above posts.

OpenStudy (anonymous):

But the way they want you to do is do it the way I put up above. Because you will get the right answer if you follow the steps and it is less complicated.

OpenStudy (inkyvoyd):

No, the way they WANT you to do it is solve the system.

OpenStudy (anonymous):

Exactly.I am prettysure I know what Iam doing. i am taking algebra II